# SiamGPT: Quality-First Fine-Tuning for Stable Thai Text Generation

### v1

@@ 6.1 Conclusion
This report introduced SiamGPT-32B, a fine-tuned variant of Qwen3-32B designed to solve a production-critical failure mode for Thai: multilingual interference that manifests as mixed-script or code-switching artifacts in otherwise Thai responses. Our core claim is that a minimal, curated supervised fine-tuning recipe can produce a reliable Thai response generator when the data is deliberately selected for control and stability rather than sheer scale. We implement this via a Quality-First translated instruction corpus paired with Thai-adapted AutoIF constraints dong2024autoif, yielding a compact but high-leverage training set.
Empirically, SiamGPT-32B shows consistent improvements over the Qwen3-32B baseline across the pillars most relevant to agentic deployments: (i) improved stability on a code-switching test (87.77 $\rightarrow$ 90.40), (ii) stronger instruction adherence on IF-Eval zhou2023ifeval (75.47 $\rightarrow$ 83.00), and (iii) substantially better multi-turn dialog control on MT-Bench zheng2023mtbench (57.94 $\rightarrow$ 75.81). On the SEA-HELM Thai leaderboard susanto2025seahelm; seahelmleaderboard, SiamGPT achieves the highest overall score among open-weight Thai models in the 30B-32B class, while also achieving strong instruction-following and multi-turn performance.
However, the SEA-HELM breakdown also highlights an important tradeoff. Despite the strongest overall score, SIAMGPT-32B underperforms compared to leading peers in the NLG (Translation) category susanto2025seahelm, plausibly because our supervision emphasizes controllability (instruction adherence and stability) and because translation-based training can introduce translationese artifacts that affect native generation scores zhang2019translationese.
@@ 6.2 Future Work
(3) Close the NLG gap without sacrificing stability: Given the observed SEA-HELM NLG weakness susanto2025seahelm, we will explicitly target Thai naturalness and generation quality while retaining deterministic constraint enforcement. Practically, this suggests adding Thai-native conversational and writing supervision and testing multi-objective training that treats stability as a hard constraint rather than a soft preference.
(4) Make corpus decisions evidence-driven via ablations: Our internal corpus experiments indicate that simply increasing data breadth degraded stability, and that DPO improved instruction-following but regressed other pillars rafailov2023dpo. In the next version, we will formalize these findings with publishable ablation tables and expand them into controlled studies (e.g., which data sources or objectives help NLG the most while preserving stability).

### v3

@@ 6.1 Conclusion
This report introduced SiamGPT-32B, a fine-tuned variant of Qwen3-32B designed to solve a production-critical failure mode for Thai: multilingual interference that manifests as mixed-script or code-switching artifacts in otherwise Thai responses. Our core claim is that a minimal, curated supervised fine-tuning recipe can produce a reliable Thai response generator when the data is deliberately selected for control and stability rather than sheer scale. We implement this via a Quality-First high-quality instruction corpus paired with Thai-adapted AutoIF constraints Dong et al. (2024), yielding a compact but high-leverage training set.
Empirically, SiamGPT-32B shows consistent improvements over the Qwen3-32B baseline across the pillars most relevant to agentic deployments: (i) improved stability on a code-switching test (87.70 $\rightarrow$ 90.40), (ii) stronger instruction adherence on IF-Eval Zhou et al. (2023) (75.47 $\rightarrow$ 83.00), and (iii) substantially better multi-turn dialog control on MT-Bench Zheng et al. (2023) (57.94 $\rightarrow$ 75.81). On the SEA-HELM Thai leaderboard Susanto et al. (2025); SEA-HELM Team (2025), SiamGPT achieves the highest overall score among open-weight Thai models in the 30B-32B class, while also achieving strong instruction-following and multi-turn performance.
However, the SEA-HELM breakdown also highlights an important tradeoff. Despite the strongest overall score, SIAMGPT-32B underperforms compared to leading peers in the NLG (Translation) category Susanto et al. (2025), plausibly because our supervision emphasizes controllability (instruction adherence and stability) which may negatively impact native natural language generation scores.
@@ 6.2 Future Work
(3) Close the NLG gap without sacrificing stability: Given the observed SEA-HELM NLG weakness Susanto et al. (2025), we will explicitly target Thai naturalness and generation quality while retaining deterministic constraint enforcement. Practically, this suggests adding Thai-native conversational and writing supervision and testing multi-objective training that treats stability as a hard constraint rather than a soft preference.
(4) Make corpus decisions evidence-driven via ablations: Our internal corpus experiments indicate that simply increasing data breadth degraded stability, and that DPO improved instruction-following but regressed other pillars Rafailov et al. (2023). In the next version, we will formalize these findings with publishable ablation tables and expand them into controlled studies (e.g., which data sources or objectives help NLG the most while preserving stability).